OpenStudy (anonymous):

x^3+k^2x^2-2kx-16 determinate k so the polynom is devidable with x-2

OpenStudy (anonymous):

if it is dividable by x -2 it should be its factor so subs x =2 in eq

OpenStudy (anonymous):

equate it to 0

OpenStudy (anonymous):

u will get k

OpenStudy (cwrw238):

if divisible by 2 2^3 + 4k^2 - 4k - 16 = 0

OpenStudy (cwrw238):

4k^2 - 4k - 8 = 0 k^2 - k - 2 = 0 solve for k

OpenStudy (cwrw238):

can you solve this?

OpenStudy (anonymous):

k=2 and k=-1?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thanks! =)